OpenStudy (anonymous):

m=4and the y-intercept is 3

OpenStudy (anonymous):

y=4x+3

OpenStudy (anonymous):

If you are looking for slope intercept form it's always y=mx+b where m is the slope and b is the y-intercept

OpenStudy (anonymous):

m=7/2and the y-intercept is -5

OpenStudy (andrew314):

You should do your own work instead of getting us to do it for you

OpenStudy (anonymous):

m=7/2and the y-intercept is -5

OpenStudy (anonymous):

use the form I showed you above^ and just put in the numbers

OpenStudy (anonymous):

[-5,7] and [0,10],,,,,,,

OpenStudy (anonymous):

substitute

OpenStudy (andrew314):

I would be more inclined to answer your question if you created a new one, so I could get points. However, I will answer your question here: (0,10) is a y-intercept point. Why? because x is 0. Then you can take (y1 - y2)/(x1 - x2) to find your slope. Use the points you are given to find the slope.

OpenStudy (anonymous):

x=3y+6